Human blood is grouped into four types. The percentages of Americans with each type are listed below.
O 43% A 40% B 12% AB 5%
Choose one American at random. Find the probability that this person
i) Has type O blood
ii) Has type A or B
iii) Does not have type O or A

Answer by ewatrrr(24785)

Choose one American at random
i) Has type O blood: p = 43/100 or .43
ii) Has type A or B: p = 52/100 0r .52
iii) Does not have type O or A = 17/100 0r .17